40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Oct 2 02:33:32 2011, kiwamu, Update, LSC, something funny with AS55 outofloop.png
    Reply  Thu Oct 6 05:18:20 2011, kiwamu, Update, LSC, Noise in AS55 was from clipping : fixed  outofloop.png
Message ID: 5595     Entry time: Sun Oct 2 02:33:32 2011     Reply to this: 5624
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: something funny with AS55 

Just a quick report.

The AS55 signal contains more noise than the REFL signals.

Why ? Is this the reason of the instability in PRMI ?

outofloop.png

 I locked the Power-Precycled ITMY with REFL33.

As shown in the plot above, I compared the in-loop signal (REFL33) and out-of-loop signals (REFL11 and AS55).

All the signals are calibrated into the displacements of the PR-ITMY cavity by injecting a calibration peak at 283 Hz through the actuator of PRM.

AS55 (blue curve) showed a structure around 3 Hz and higher flat noise below 1 Hz.

Quote from #5582
I am suspecting that something funny (or stupid) is going on with the MICH control rather than the PRCL control.

 

ELOG V3.1.3-